### v1

@@ Abstract
Real-world development demands code that is readable, extensible, and testable by organizing the implementation into modular components and iteratively reuse pre-implemented code. We term this iterative, multi-turn process codeflow and introduce CodeFlowBench, the first benchmark designed for comprehensively evaluating LLMs’ ability to perform codeflow, namely to implement new functionality by reusing existing functions over multiple turns. CodeFlowBench comprises 5,258 problems drawn from Codeforces and is continuously updated via an automated pipeline that decomposes each problem into a series of function-level subproblems based on its dependency tree and each subproblem is paired with unit tests. We further propose a novel evaluation framework with tasks and metrics tailored to multi-turn code reuse to assess model performance. In experiments across various LLMs under both multi-turn and single-turn patterns. We observe models’ poor performance on CodeFlowBench, with a substantial performance drop in the iterative codeflow scenario. For instance, o1-mini achieves a pass@1 of 20.8% in multi-turn pattern versus 37.8% in single-turn pattern. Further analysis shows that different models excel at different dependency depths, yet all struggle to correctly solve structurally complex problems, highlighting challenges for current LLMs to serve as code generation tools when performing codeflow. Overall, CodeFlowBench offers a comprehensive benchmark and new insights into LLM capabilities for multi-turn, iterative code generation, guiding future advances in code generation tasks. The code and dataset will be available soon on https://github.com/IAAR-Shanghai/CodeFlowBench
@@ 1 Introduction
Large Language Models (LLMs) have been widely applied to code generation tasks and many benchmarks have been developed to evaluate LLMs’ performance on code generation tasks, such as HumanEval[3], MBPP[2], and DS-1000[15]. With the rapid advancement of LLMs’ capabilities, the focus of using LLMs for code generation has shifted from merely tackling toy problems to integrating these models into real-world development workflows for more complex code implementations with the rapid advancements of the capabilities of LLMs.Correspondingly, several benchmarks have begun to ground their evaluation content in practical development scenarios. For instance, BigCodeBench[29] selects functions directly from open-source repositories and requires models to invoke external libraries while SWE-Bench[13] evaluates bug-fixing capabilities via GitHub pull requests.
However, we contend that current benchmarks still fail to capture two fundamental and pivotal aspects of real-world development workflows: multi-turn and iterative. To enhance the readability, extensibility and testability of code, developers tend to organize their workflow as a codeflow where they start by implementing the lowest-level functions to solve fundamental subproblems and then reuse these functions to build more sophisticated modules. An example of codeflow is shown in Figure 1. Codeflow is inherently a multi-turn, iterative code generation process since developers refine or extend functionality by leveraging pre-implemented code in each turn.
Figure 1: An example of building a currency arbitrage detection and execution system. The implementation in left side is organized in a codeflow format where the developer first creates three modular functions and then reuses them in the main function to process input data and determine optimal strategies. This approach allows for easier testing, debugging and extension of each individual component. In contrast, the right side shows a single-turn implementation where the entire process is combined into one block, making it harder to read, test and maintain.
To serve as an excellent code-generation tools in a codeflow, LLMs must handle developers’ multi-turn requests, understand the dependency relationships between existing code and newly requested code in each turn and then produce reusable code for subsequent iterations. However, we find that existing benchmarks like MTPB[21] and CodeEvalPro[28] lack sufficient difficulty and do not provide a comprehensive, fine-grained evaluation framework for LLMs’ multi-turn, iterative code generation capabilities. Therefore, we believe there remains a need for a more challenging benchmark that can thoroughly evaluate LLMs’ capabilities in a codeflow.
To fill the gap, we introduce a new evaluation benchmark called CodeFlowBench. Unlike benchmarks mentioned above, CodeFlowBench sources its problems from the competitive programming platform Codeforces [^1] https://codeforces.com/, which offers several advantages:(1) High difficulty. Even state-of-the-art LLMs have low pass rates on Codeforces problems[22]. Compared to simple and straightforward code implementation, competitive problems can make multi-turn, iterative solution process more realistic and meaningful. (2) High quality. Unlike many open-source repositories, Codeforces publishes official editorial and provides automated correctness checks, ensuring the integrity of our dataset’s problems.(3) Real-time updates. To maintain uncontaminated test data and ensure fairness, it is essential that benchmarks can be refreshed regularly. Codeforces releases new problems at a high frequency, enabling continuous updates for CodeFlowBench.We also developed an automated, lightweight pipeline that processing problems in Codeforces to be sets of multi-turn, iterative subproblems to build CodeFlowBench and make it available for continuous updates.
@@
In this paper, we introduce the initial version of CodeFlowBench, which comprises 5,258 test problems extracted from the Codeforces problem archive accumulated until now. To ensure ongoing fairness and freshness, we will periodically run our automated pipeline and continuously update the benchmark with new, uncontaminated problems.
In sum, the contribution of this paper is threefold:
@@
  • We build a data curation pipeline based on the Codeforces platform. This pipeline decomposes existing solutions based on function dependencies to form multi-turn, iterative code problems that require code reuse. The approach is fully automated, lightweight, and easy to maintain, allowing for real-time updates.
@@
  • We introduce the first benchmark to specifically target the multi-turn, iterative code generation capabilities of LLMs. It features a large and diverse collection of competition-level code problems. By leveraging official editorial solutions, automatically verifying generated code and capturing test cases through input-output streams, we ensure the robustness and quality of our dataset.
@@
  • We propose a novel evaluation task framework that incorporates both multi-turn and single-turn code generation patterns. We also introduce new structural labels and metrics derived from the dependency tree of each problem, aiming to more comprehensively capture the characteristics of multi-turn tasks and bridge the gap between existing evaluation metrics and the unique challenges posed by multi-turn tasks. The experimental findings reveal important characteristics of model capabilities and deficiencies in multi-turn, iterative code generation, offering meaningful directions for future research and development.
@@ 2 Related Work
The landscape of code generation benchmarks has been progressing from simpler to more complex tasks. Pioneering works like HumanEval[3] and MBPP[2] center on code generation for stand-alone function. The main constraints are simple problem difficulty and limited dependency environments. Building on these works, recent benchmarks have emerged to evaluate various aspects of LLMs as code generation tools: Benchmarks such as APPS[7], CodeContest[17], LiveCodeBench[11], and CodeElo[22] employ complex competition-style problems to rigorously test the models’ ability to understand and reason about code; Benchmarks like MTPB[21], InterCode[27] and TreeInstruct[14] focus on the iterative, multi-turn dialogue approach users often adopt when generating code with LLMs, thereby evaluating the models’ ability to retain context, understand, and reason across multiple interactions. Meanwhile, CodeEvalPro[28] uses existing functions to create self-invoking functions, testing the models’ ability to solve these tasks by invoking existing code. Besides, LiveCodeBench[11] and EvoCodeBench[16] plan regular updates to their problem sets, aiming to mitigate the risk of data leakage and maintain testing fairness. Collectively, these benchmarks significantly contribute to the development and refinement of evaluation frameworks. Our work builds upon these cutting-edge benchmarks, focusing on the multi-turn, iterative code generation needs of developers tackling complex implementations, and offering a novel, high-quality benchmark to address these challenges.
@@
In recent years, large language models have achieved remarkable and rapid progress in code generation. Early works such as Codex[3] and AlphaCode[17] have demonstrated impressive performance in a wide range of tasks from simple function completion to solving competition-level programming problems. With the scaling up of pre-trained models, exemplified by GPT-4[1], Code-Llama[23], Deepseek-Coder[5], and Qwen2.5-Coder[9], advanced pre-trained models now have high-level performance across various programming tasks in different languages and domains. Building on these foundations, the code generation capabilities of LLMs have further advanced through instruction tuning and agent frameworks. Instruction tuning Models like WizardCoder[19] and MagicCoder[25] exhibit stronger interactive dialogue capabilities and a better grasp of user intent, while agent-based frameworks such as Agentcoder[8] and Mapcoder[10] enhance the autonomy of LLMs in code generation and enable iterative planning, execution, and self-evaluation in code generation task. Overall, LLMs are rapidly becoming indispensable tools for developers, significantly reducing the time and effort required to produce high-quality code.[12]. Our work provide a new evaluation approach and direction for the advancement of LLMs’ code generation capabilities by targeting the multi-turn, iterative feature of code generation.
@@ 3 CodeFlowBench
CodeFlowBench is a lively-updated benchmark which contains 5,258 diverse, high-quality and challenging problems.An example problem of CodeFlowBench is shown in Figure 2. Remaining part will introduce the data curation pipeline and the evaluation method of CodeFlowBench.
Figure 2: An example problem from CodeFlowBench, which is originally sourced from problem 2055E in Codeforces.The top-left block displays the whole problem information, including scraped data as well as the designated overall-depth and overall-turns metrics. The bottom-left block illustrates the dependency structure of the problem, where each node represents a subproblem and each connecting line indicates a dependency relationship. The right block details each subproblem. Models are required to engage in this codeflow by solving the subproblems iteratively, turn by turn. The bottom block shows the LLM interaction workflow for each turn.
@@ 3.1 Data Curation Pipeline
We developed an automated, lightweight data-curation pipeline for generating multi-turn coding problems. The pipeline consists of two main stages: (1)Scraping and processing data (Stages I–III): it contains source-specific routines that collect and normalize raw input from various repositories. (2)Decomposing and generating subproblems (Stages IV–VI): it contain a general framework that, treats each function as an individual subproblem and decomposes the problem according to the functions’ dependency relationships given a codebase.
It is the first pipeline to incorporate the official editorials of Codeforces problems and to include a submission bot for fair verification across all Codeforces-targeted pipelines. By incorporating the official editorials, we avoid scraping user solutions which may raise ethical concern, eliminating the need for extensive filtering and ensure that the codeflow for each problem is organized within an optimal framework, thereby preserving the high quality of our benchmark. We believe our pipeline can also contribute to future competitive code generation benchmarks. Specifically, our pipeline comprises the following steps.
Figure 3: Illustration of the data curation pipeline of CodeFlowBench
@@
We visit the Codeforces problem list [^2] https://codeforces.com/problemset and then navigate to each problem’s page. We extract the following information: problem ID, title, time and memory complexity limits, problem description, input description, output description, sample tests, notes. Additionally, by querying the official Codeforces API, we retrieve each problem’s rating and tags. The detailed explanation of the information and an example can be found in Appendix A.1.
@@
We visit the tutorial page from problem website. Since tutorials are typically collections of problems grouped by a contest round, we opted to scrape explanations at the round level. The explanation pages display each problem’s identifier, textual explanation and code solution in sequence. Therefore, we first locate each problem’s identifier and extracts the content between adjacent identifiers as the solution content of corresponding problem. We primarily rely on the URL corresponding to each problem ID in the HTML code for identifier recognition, supplemented by common textual patterns to enhance accuracy. Once the scraping is complete, we map these identifiers to the problem IDs collected in Stage I to append the explanations to each problem. Additionally, we filter out explanations of evidently low quality and split the explanations into text and code segments to facilitate model processing. A detailed explanation of the method and an example can be found in Appendix A.2
@@
We feed both the problem and its explanation information into the Deepseek-R1 model[6], allowing it to either directly use the code implementation if provided or organize the existing textual solution and convert it to parsable solution code that meets the requirements for extracting function dependencies and constructing subproblems. The generated code is then submitted to the Codeforces platform for correctness verification, and we retain only those codes that successfully solve the problem. The prompt template can be found in Appendix A.3
Stage IV: Subproblems Generation
We first parsing code generated in Stage III with its Abstract Syntax Tree(AST) to obtain its function dependency relationship. To select potential subproblem and correctly arrange it as a codeflow, we filter out low-quality functions and applying topological sort based on dependencies to ensure that lower-level functions precede higher-level ones. Each filtered function is treated as a subproblem. For every function, we extract its corresponding code segment from the solution and record its depth in the AST along with the functions it depends on. We then input this information, together with the problem and explanation details, into the Deepseek-V3 model[18] and require the model generates a problem description for each function implementation, thereby forming a complete subproblem that contain name, statement, depth, dependencies, solutions attributes.The prompt template and an example can be found in Appendix A.4
@@
For each problem, we first scrape the publicly available test cases from the Codeforces platform and use these as inputs to execute the corresponding solution code. Meanwhile, we embed input-output capture logic at every location where a function is invoked within the solution code. During execution, we record the input-output pair for each function call. These input-output pairs serve as test cases for each subproblem. We explain the process in detail in Appendix A.5.
@@
For each problem, we define two metrics as labels to indicate the complexity of the problem’s structure, which are useful for our experiment analysis.
  • Overall-Turns: The number of turns required to solve the entire problem.
  • Overall-Depth: The depth of the AST. Each subproblem serves as a node in the tree, with the main/solve function as the root node.
The distribution of overall-turns and overall-depths in CodeFlowBench is shown in 4(a). We also compare the Rating metric provided by Codeforces platform which reflects the difficulty of a problem with our metrics.The strong correlation shown in 4(b) confirms that our label effectively captures the intrinsic complexity of these problems.
Figure 4: Statistics of the overall-turns and overall-depth metrics in CodeFlowBench. Subfigure (a) illustrates the distribution of these two metrics. Subfigure (b) depicts their correlation with the Rating metric.
@@ 3.2 Evaluation Framework
Partially like the task defined in HumanEval[3], models evaluated by CodeFlowBench are also required to complete the code implementation for a given function in each turn. The main difference lies in the supporting materials of the task. In addition to the function signature and problem description, we also provide dependency information and pre-implemented functions. We require models to reuse existing code if needed. Moreover, we provide an overall background description of the whole problem since we consider the opportunistic scenario where developers have knowledge of the whole problem as a ultimate goal when implementing code in a turn to be more common and reflective of real-world situations.
The evaluation task can be formally defined. Suppose $F_{t}$ is the function for which the implementation $C_{t}$ is to be generated at turn $t$ and $S_{t}$ is the corresponding problem statement for $F_{t}$, which may include the description of inputs, outputs, and functionality of $F_{t}$. $\mathcal{D}=\{k_{1},k_{2},\ldots,k_{d}\}$ denote the set of previous turns on which the current turn depends and $F_{k_{1}},F_{k_{2}},\ldots,F_{k_{d}}$ represent the definitions of dependency function. $B$ is the background information for the overall problem and $C_{1},C_{2},\ldots,C_{t-1}$ denote the implementations generated in the chatting history. Suppose the total number of turns is $T$. Then the implementation $C_{t}$ at turn $t$ is generated as follows $(t=1,2,\cdots,T)$:
$C_{t}\leftarrow p\Bigl{(}\cdot\,\Bigm{|}\,F_{t},S_{t},\{F_{k_{1}},F_{k_{2}},\ldots,F_{k_{d}}\},\{C_{1},C_{2},\ldots,C_{t-1}\},B\Bigr{)}.$
For comparison, we also provide a single-turn evaluation pattern to assess the model’s multi-turn capabilities. In the single-turn task, the final implementation is generated by:
$\{C_{1},C_{2},\cdots,C_{T}\}\leftarrow p\Bigl{(}\cdot\,\Bigm{|}\,\{F_{1},F_{2},\cdots,F_{T}\},\{S_{1},S_{2},\cdots,S_{T}\},B\Bigr{)}.$
Pass@k [3] is a widely used evaluation metric for code generation tasks. We adopt it as our primary evaluation metric for both multi-turn and single-turn evaluation patterns. However, a notable limitation of using $pass@k$ in a multi-turn setting is that it only provides a final, coarse-grained evaluation and cannot fully distinguish the capabilities of different models. In other words, pass@k only evaluate models’ performance on a problem with subproblems based on whether the model pass the whole problem.Such method will evidently loss many information since different models may have different performance on subproblems even though they all fail to solve the whole problem.
Therefore, it’s much better to locate the exact turn $t$ at which the model fails by using unit tests for each subproblem. However, directly comparing failing turns may be unfair because the topological ordering of subproblems at same depth in AST can be arbitrary. To address this issue, we propose the Pass Depth (PD) metric. For a problem with overall-depth $D$, we define its PD as follows:
$PD=\begin{cases}D-d-1,&\text{if the model makes an error on a depth-$d$ subproblem;}\\ D,&\text{if the model passes the whole problem.}\end{cases}$
We will report the Average Pass Depth (APD) by averaging the PD values across problems grouped by depth. In the case of $k$ trials, we define PD(k) as the maximum PD achieved over $k$ attempts, and we use PD(k) to compute the corresponding APD(k).

### v2

@@ Abstract
Modern software development demands code that is maintainable, testable, and scalable by organizing the implementation into modular components with iterative reuse of existing codes. We formalize this iterative, multi-turn paradigm as codeflow and introduce CodeFlowBench, the first benchmark designed to comprehensively evaluate LLMs’ ability to perform codeflow – implementing new functionality by reusing existing functions over multiple turns. CodeFlowBench comprises 5,258 problems from Codeforces and is continuously updated via an automated pipeline, which decomposes each problem into subproblems with unit tests based on dependency tree analysis and dataflow analysis. We further propose a novel evaluation framework featured dual assessment protocol and structural metrics derived from dependency trees. Extensive experiments on 16 popular LLMs reveal significant performance degradation in multi-turn scenarios. For instance, o1-mini retains only 20.8% Pass@1 in multi-turn scenario versus 37.8% in single-turn scenario. More fine-grained analysis illustrates that model performance inversely correlates with dependency complexity. These findings not only highlight the critical challenges for supporting real-world workflows, but also establish CodeFlowBench as an essential tool for advancing code generation research.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized code generation, with benchmarks like HumanEval [5] and MBPP [3] establishing foundational standards. As LLM capabilities advance, their role in real-world software development has expanded beyond solving toy problems to supporting complex workflows [21, 23, 28]. Modern benchmarks such as DevBench [25] and SWE-Bench [22] now emphasize practical scenarios like bug fixing. However, current benchmarks [9, 15, 35] still overlook the critical aspect of real-world development: the multi-turn and iterative codeflow scenario.
The CodeFlow Task
In modern software engineering, multi-turn and iterative workflows are becoming increasingly prevalent, as the cornerstones of best practices like agile development [24, 1, 6]. By breaking down complex tasks into manageable subproblems, progressively refine solutions and reuse modular functions, developers can achieve faster delivery, reduced redundancy and enhanced maintainability in teamwork [12, 7]. For example, React’s core package alone sees over 37 million weekly downloads across 2 300+ dependent modules, illustrating the productivity gains of modular reuse [4]. As shown in Fig. 1 and 2, by building solutions by traversing the function dependency tree from the bottom up, the structured codeflow can both enables parallel development and improves readability, testability and maintainability through modular and responsibility boundaries. For LLMs to integrate effectively into such workflows, they must handle multi-turn interactions, manage dependencies between code segments and generate reusable components for subsequent iterations.
Figure 1: An example of building a currency arbitrage detection and execution system. The left side demonstrates a codeflow workflow, where modular functions are reused to construct the main logic. The right side represents a single-turn implementation, combining all steps into a monolithic block.
Motivation
Despite the growing demands, current benchmarks have not fully captured the multi-turn and iterative aspects of codeflow. Firstly, most benchmarks, such as HumanEval and MBPP, focus only on single-turn code generation. While recent benchmarks like BigCodeBench [40] and SWE-Bench [22] have begun to incorporate practical development scenarios, they still remain at the level of single-turn code modification (or two-turn code generation), leaving the multi-turn generation capabilities unknown. Secondly, the few existing multi-turn benchmarks such as MTPB [31], only focus on single-function programming, lacking both unit tests and sufficient complexity to reflect real-world dependencies. Finally, due to the absence of an update mechanism, previous static datasets risk contamination and unreliable assessment. Therefore, there is a pressing need for a more challenging, well-annotated and frequently updated benchmark specially designed for codeflow.
To bridge this gap, we propose CodeFlowBench, a novel evaluation benchmark that sources its problems from the competitive programming platforms like Codeforces [^1] https://codeforces.com/, featuring high difficulty (competition-level problems), high quality (peer-reviewed editorial solutions), and periodic updates (frequent problem additions), elaborated in Appendix A. CodeFlowBench leverages platforms’ frequent release of new problems to ensure continuous updates for the uncontaminated benchmark. Based on dependency decomposition and online submission check, an automated, lightweight pipeline is further developed to process original problems into sets of multi-turn, iterative subproblems with verified test cases and solutions. Remarkably, CodeFlowBench also introduces a specialized evaluation framework, including structural labels and metrics derived from dependency trees, to comprehensively assess multi-turn performance and provide more insights for model shortcomings.
Contributions
@@
In this paper, we present the first edition of CodeFlowBench, which comprises 5,258 complex problems fetched from the Codeforces problem archive accumulated until now. To ensure ongoing fairness and freshness, we will periodically perform the automated pipeline and update the benchmark with brand-new, uncontaminated problems. Our main contributions are threefold:
@@
  • Pipeline Innovation: We develop a data curation pipeline based on competition platform. This pipeline decomposes official solutions based on function dependencies to create multi-turn, iterative code problems that require code reuse. This lightweight pipeline is fully automated, rigorously verifiable, and easy to maintain, facilitating frequent updates.
@@
  • Benchmark Construction: We introduce the first benchmark specifically designed to evaluate the multi-turn, iterative code generation capabilities for codeflow settings. It features a large and diverse collection of 5,258 complex code problems. By leveraging official editorial solutions, automatically verifying generated code, and capturing test cases through input-output streams, we ensure the quantity and quality of our dataset.
@@
  • Evaluation Design & Insights: We propose a novel evaluation framework that contrasts multi-turn and single-turn code generation patterns. Our framework introduces new structural labels and metrics derived from the dependency trees, aiming to better capture the unique characteristics of multi-turn tasks. These innovations bridge the gap between traditional evaluation metrics and the specific challenges posed by iterative code generation. Extensive experiments reveal the huge performance drop in the codeflow scenarios, even for large reasoning models, highlighting the need for more advanced multi-turn capabilities.
@@ 2 Related Work
The landscape of code generation benchmarks has evolved from simpler to more complex tasks to keep pace with the rapid development of LLMs [26, 32, 39], but still fail to comprehensively capture the multi-turn and iterative features of real-world scenarios. Early works like HumanEval [5] and MBPP [3] focus on standalone functions with low complexity and limited dependency environments. Recent benchmarks have emerged to evaluate more complex and realistic scenarios, yet have obvious limitations: most benchmarks such as APPS [13], LiveCodeBench [20] and SWE-Bench [22] are limited to single-turn code generation or modification. For the few existing multi-turn benchmarks, MTPB [31] focus on overly simplistic single-function programming without paired unit tests, while InterCode [38] discusses interactive coding with execution feedback. In stark contrast, codeflow structures the development into multi-turn processes, ensuring each component is maintainable, testable and reusable. These limitations highlight the crucial gaps for codeflow benchmarking, and the CodeFlowBench pioneers this research line.
@@
Recent years have witnessed unprecedented progress in code generation capabilities of LLMs. Early works such as Codex [5] and AlphaCode [26] demonstrated proficiency in tasks ranging from code completion to competition-level problem solving. With the scaling up of pre-trained models, exemplified by GPT-4 [2], Code-Llama [33], Deepseek-Coder [10] and Qwen2.5-Coder [16], these advanced models have impressive performance across various programming tasks, languages and domains. Building on these foundations, the code generation capabilities have further advanced through instruction tuning and agent frameworks. Models such as WizardCoder [29] and Magicoder [36] leverage instruction tuning to improve intent alignment and interactive dialogue capabilities, while agent frameworks like AgentCoder [14] and MapCoder [18] enable autonomous planning, iterative refinement, and self-evaluation. Despite these advancements, the community still remains unknown about "how well and how deeply LLMs can perform codeflow"—a critical paradigm for real-world software engineering. Our CodeFlowBench thus provides a principled framework for advancing both model development and evaluation [21].
@@ 3 CodeFlowBench
CodeFlowBench is a lively-updated benchmark which currently contains 5,258 diverse, high-quality and challenging problems. An example problem of CodeFlowBench is shown in Figure 2. In this section, we introduce the data curation pipeline and evaluation framework of CodeFlowBench.
Figure 2: An example problem from CodeFlowBench, sourced from Codeforces problem 2055E. The top-left block displays the problem information. The bottom-left block illustrates the dependency structure, with nodes as subproblems and edges as dependency relationships. The right block details each subproblem, where models solve these subproblems iteratively. The bottom block shows the LLM interaction workflow for each turn.
@@ 3.1 Data Curation Pipeline
We designed an automated, lightweight data-curation pipeline to generate complex multi-turn coding problems. The pipeline consists of two main phases: (1) Data Preparation (Stages I–III), which involves source-specific routines tailored to collect and normalize raw problem data from various platforms. (2) Subproblem Generation (Stages IV–VI), which employs an universally applicable problem decompostion framework. This framework systematically analyzes dependency relationships between functions or classes, and back-translates functions into coherent and sequential subproblems. Notably, our implementation is the first to integrate both official problem editorials and automated submission checker into the data curation process. Leveraging official editorials and submission checker provides reliability benefits, such as eliminated copyright issues or data cleaning needs, and it also ensures the codeflow implementation is high-quality and optimal for each problem. As shown in Fig. 3, our pipeline specifically comprises the following stages:
Figure 3: The data curation pipeline. Left: in the data-collection stage, we scrape problems and corresponding text and code solutions. Middle: in the data-processing and subproblem-generation stage, we validate solutions, parse codes into abstract syntax trees, decompose original problems into individual subproblems, and generate test cases. Right: the final labeling stage for complexity analysis, and the curated CodeFlowBench problem set.
@@
We initiate the process by visiting the Codeforces problem list and navigating to each problem’s page to extract the following information: problem ID, title, time and memory limits, problem description, input/output specifications, sample tests, notes. Furthermore, by querying the official Codeforces API, we retrieve each problem’s rating and tags. Please refer to Appendix B.1 for information details and examples.
@@
We gather editorial content from corresponding tutorial pages, where each problem’s ID, textual explanation, and solution code are listed sequentially. By identifying problem IDs through URL and text patterns, we extract the relevant content and associate it with the problem list generated in Stage I. Low-quality entries are filtered out, and solutions are categorized into text and code components. Please refer to Appendix B.2 for more details.
@@
We input both the problem and its explanation information into the Deepseek-R1 model [11], which either directly uses the provided code implementation or organizes the existing textual solution and converts it into parsable solution code that meets the requirements for extracting function dependencies and constructing subproblems. The generated code is then submitted to the Codeforces platform for correctness verification, and we retain only those codes that successfully solve the problem. The prompt template can be found in Appendix B.3.
Stage IV: Subproblem Generation
We parse the abstract syntax tree (AST) of codes from Stage III to extract function dependencies, filter out built-in functions, and conduct topological sort so that lower-level ones come first. We then treat each function as a subproblem and record its solution code, AST depth, and dependencies information. These information, along with the problem and explanation, are fed into Deepseek-V3 [27] to back-translate a subproblem description. Finally, a subproblem contains the following fields: name, statement, depth, dependencies, solution. Detailed explanations of the prompt template and an example can be found in Appendix B.4.
@@
For every problem, publicly accessible test cases from the Codeforces platform are scraped and used to execute the respective solution codes. During execution, we capture input-output pairs at every function call location, which serve as test cases for the corresponding subproblems. Further elaboration on this process can be found in Appendix B.5.
@@
To quantify the complexity of the curated problems for fine-grained analysis, we define two metrics: Overall-Turns–the number of turns required to solve the entire problem; Overall-Depth–the depth of the AST, where the main/solve function as the root and subproblems as nodes. As shown in Fig. 4, we report the distributions of overall-turns and overall-depths, and correlate them with the difficulty ratings provided by Codeforces. The strong correlation observed confirms that our label effectively captures problems’ intrinsic complexity.
(a) Distributions of overall-turns and overall-depth.
(b) The Correlations with Rating Levels.
Figure 4: Statistics of the overall-turns and overall-depth metrics in CodeFlowBench. Subfigure (b) shows inflection points at turns = 1 and depth = 1. This is attributed to the fact that competition-level problems are not restricted to multi-turn or deeply nested structures. Partial difficult problems are designed to be solvable by single function such as number theory related problems.
@@ 3.2 Evaluation Framework
Task Definition
The evaluation design of CodeFlowBench follows best practices defined in HumanEval [5] but introduces a novel multi-turn iterative paradigm. That is, models are required to implement a given function in each round. However, CodeFlowBench introduces distinct differences in the supporting materials provided to the models. Beyond the function signature and problem description, we include dependency information and pre-implemented functions for code reuse. Additionally, we provide an overarching background description of the problem. This reflects the real-world scenario where developers often possess overall understanding of the entire problem when implementing code incrementally.
As illustrated in Fig. 2, models must implement target functions while leveraging dependency relationships and pre-implemented components from prior turns. Formally, given $T$ total turns, the function implementation $C_{t}$ at turn $t$ is generated based on:
  • Function signature $F_{t}$ and subproblem description $S_{t}$ for the current target function,
  • Background context $B$ describing the overarching programming problem,
  • Dependency specifications $\mathcal{D}=\{k_{1},...,k_{d}\}$ indicating prerequisite functions $F_{k_{1}},...,F_{k_{d}}$,
  • Verified functional implementations $\{C_{1},...,C_{t-1}\}$ of previous dependencies.
For baseline comparison, we define the single-turn variant where the model generates all components simultaneously. The mathematical expressions of both settings are presented below:
$C_{t}\leftarrow p$ $\Bigl{(}\cdot\,\Bigm{|}\,F_{t},S_{t},\{F_{k_{1}},F_{k_{2}},\ldots,F_{k_{d}}\},\{C_{1},C_{2},\ldots,C_{t-1}\},B\Bigr{)}\quad(1\leq t\leq T),$
$\{C_{1},C_{2},\ldots,C_{T}\}\leftarrow p\Bigl{(}\cdot\,\Bigm{|}\,\{F_{1},F_{2},\ldots,F_{T}\},\{S_{1},S_{2},\ldots,S_{T}\},B\Bigr{)}.$
Performance Metrics
We adopt widely-used Pass@k [5] as the main metric for both multi-turn and single-turn cases, but augment it with novel diagnostics for multi-turn analysis. While Pass@k evaluates final success rates, the coarse-grained nature fails to capture partial progress in failed attempts - two models may fail solve the entire problem at different stages but receive identical scores.
To address this limitation, we propose a new metric, Pass Turn ($PT$), which identifies the exact turn at which a model fails by leveraging unit tests for each subproblem. However, only considering failing turns may be biased due to the arbitrary topological ordering at the same depth. We further define, Pass Depth ($PD=D-d$), for bottom-up programming, where $d$ and $D$ are working depth and total depth. For statistical significance, we define the Average Pass Turn (APT) and Average Pass Depth (APD) by averaging across problems grouped by turn or depth. For multiple trials, we define APD@k and APT@k following the Pass@k expression. Refer to Appendix C for more details.